UT 20131110 LUCI2 Nighttime Commissioning

Observer: WSeifert, JHeidt, PBuschkamp, APramskiy, KPolsterer,MEdwards, DThompson
Telescope Operator: DGHuerta
LBTO Support: TSargent, DMiller

Plan for the night

  • => First Com-night, concentrate on basics first, adapt to weather if necessary, if bad, do closed-dome stuff

  • Note1: Since most likely no closed-dome tests could be done on Nov 8, some important tests may need to be done before going on sky
  • Note1: No telescope service GUI will be available during this run, ie all has to be done via scripts
  • DMiller will be on call from 10pm onwards in case guide probe transformation measurements are possible

  • Pupil alignment verification (on sky)
  • Preset by TO to collimate and check if pointing star can be seen
  • Preset in track/active at 3 positions mode to see if it works (pointing star)
  • Redo with offsets to have a feeling of orientation
  • Orientation on sky, Pearsson stds, about 12 fields (if weather photometric, some of them in 4 filters, can be usd for first guess of ZPs)
  • Telescope focus, TT (if seeing is good and conditions are stable, need DMiller)
  • Guide probe transformation (if seeing is good by DMiller)
  • Take image of with pointing mask
  • Offset tests small/large, jitter-large offset sequences, failure tests (do offset with GS off patrol field)
  • Guide probe transformation verification (if seeing is good) by DMiller)

Summary

Successfull first LUCI2-com night. Lots of bugs found, some removed.
Presets and offsets via scripts work. Did quite a few ZP measurements mainly in BB-filters.
Night was photometric about 3/4 of the night, some cirrus moved in at twilight.

Telescope worked mostly well, had some problems with collimation which amount to one hour loss.

Main bugs to be fixed
  • Within a script data should be taken once telescope is collimated
  • Fits-header must be correct
  • Catalog for GS-selection must be available on site

Preparations

MEdwards has prepared quite a few scripts for pointing stars/Pearsson stds.
DThompson provided info about his procedure for closed-dome pupil alignment
Flexure-script to check flexure compensation in steps of 60degs created (flexure_check.xml)

Details (times are in UT)

Files are luci2.20131109*

23:00 Handover of telescope

23:57 Starting check of flexure compensation (#22=dark,#23/24 test images)
Script didn't start, Alexander has to check

00:00 In the meantime Date is Nov 10 2013, files are now luci2.20131110*

Check flexure compensation

00:20 Flexure check script started (had to restart data dictionary).

00:29 Sunset, clear skies, +4deg, 50% humidity, low wind

00:20 Restart flexure script (flexure comp was off before), files #39-56

Walter checked flexure compensation: it works!!!!!

Note: ALADIN automatically load images, but gives error message
Failed to load file from remote call: Exception: java.lang.nullPointerException

00:42 Instrument manager hangs => restart sequencer, OK

00:57 Point to a bright star (HD 223952), preset failed

01:01 Redone by TO, let telescope collimate, OK, tried to take image=> GEIRS problem, reset

01:10 OK, took image of star (#56), orientation on detector is wrong (known from yesterday), modify GEIRS,
FWHM is 0.44"

01:23 Took image #57, orientation is OK!!!, change was in start_LUCI_new, DET_XYflip 2

01:27 Put pupil viewer in, instr.manager got stuck, PV-image #58, looks decent, redo (#59), OK

01:38 Preset to ACT0244, got error message about rotator, since both sides are authorized

01:55 Done, restart script, problem with FW2, manually fixed

02:16 Preset to ACT0244, mirror panicked, in the meantime Alex & Kai deployed a new SW-version

02:38 Preset to ACT0244, OK, error message related to instrument-manager, restart SW

02:44 Preset to ACT0244, OK, FIRST images taken with a script (#61,62)

02:50 Orientation on sky: preset to Stone B (StoneB _st102), preset, OK, error message for instr.manager again,
David need to home the mirror

02:54 Redo preset to Stone B, OK, but script started before telescope was collimated. Takes a long time (>15min),
called DMiller, fixed it, collimated now

If a script is executed, taking data starts right away once preset was successfull, ie doesn't wait for collimation
Until this is fixed, script is started and stopped as soon as the first observing item is being executed
Than wait for collimation and restart script => these data are useful

03:39 Redo preset to Stone B, OK,. FWHM 0.55", #65+66

Orientation on sky

03:43 Preset to Stone B (StoneB _sl01), again problems with collimation, restart GCS

03:59 StoneB _sl01 collimated (#70+71)

04:03 StoneA _sl01, OK collimated (#72+73)

04:07 StoneA _sl02, collimation error, need to wait again...

04:07 StoneA _sl02, preset failed, David had to correct pointing

04:19 StoneA _sl02 barely collimated (#75+76), 0.5" FWHM

04:22 StoneA _sl03, preset failed, problem in detecting GS due to high background, go to new field

04:27 StoneC _sl01, preset failed, GS too far away?? (potentially scripting error?)

04:40 BS9188, preset failed again, now failed for 3 fields in a row, source of error is not SW-checking

04:58 Modified script for BS9188 by setting "guideStarsAuto" flag from true to false

05:00 Preset to BS9188 worked, can also for ZP in Ks be used(#79-83)

Suspect that automatic GS-search by the planning tools or telescope service is based on online-catalog,
whose server might be down from time to time

05:18 Preset to StoneC _sl01, no GS, check pointing, redo, preset again failed, GS is presumably a galaxy

05:27 Preset to StoneC _sl02, OK, #84+85

05:38 Preset to StoneC _sl03, OK, #86+87

05:43 Preset to StoneD _sl01, OK, #88+89

05:47 Preset to StoneD _sl02, OK, #90+91

05:51 Preset to StoneD _sl03, OK, #92+93

05:55 Preset to StoneD _sl04, OK, #94+95

05:59 Preset to BS9116, (also ZP) OK, 0.6" FWHM, #97-101

06:11 Preset to BS9105, (also ZP) , preset failed, GS looks crazy, David need to collimate first, OK, #102-106

Fits-header are crappy, a couple of entries are bad, eg RA is entered twice with different positions

Dave tries to find out a way to astrometrically calibrate the images, otherwise no
AGW-transform data can be taken tonight

06:26 Preset to BS9104, (also ZP), no good GS + problems with AGw, skipped

06:48 Preset to BS9101, (also ZP), OK, #110-114

Photometric Zeropoint's in JHKKs

06:59 Preset to BS9108, OK, AM=1.126, FWHM=0.6", #116-132,
did JHK OK, but failed after two KS-frames

07:32 Preset to BS9107, found GS, but failed "guiding is not active", restart GCS, OK,
Am=1.003-, #133-137 (J-filter), got stuck in H, skipped, in K-band at 2. exposure
suddenly preset was sent, stopped and let the SW-people check

10:05 Resumed BS9107 script in H, K, Ks, seeing degrades, now about 1.1", files #153-157

Offsets in scripts are in fact presets and thus pretty time-consuming.
To test whether this is a bug, a test script was made where the tag "mount" was deleted for the offsets 2-5 out of 5.

10:32 Preset to BS9107, zJ_spec, script BS9107_test, with 1 preset and 4 real offsets
=> works, ie bug in scripting tool

10:46 BS9107, ZP in Br_gam & OH_1060, #173-176, readout-manager hangs after the 4th Br_gam exp,
SW-people check, seeing improved to 0.5", restart GEIRS

11:23 BS9107, ZP in Br_gam & OH_1060, AM=1.47, crazy, gets now stuck at
3rd exposure in OH_1060, #177-181 (Br_gam), #182-183 (OH_1060),
need to restart telescope service

12:05 BS9108, present cancelled, shell ripped, too low elevation

12:09 BS9107, JHKKs, real offsets, AM=1.8, got stuck after J-band, restart telescope service

12:32 BS9107, JHKKs, real offsets, AM=2, #191-197, stopped after 3rd H-band exposure
fiddled a bit with various services, was able to restart

13:19 J-flats 10000-30000 cts, #205-244, ReadoutManager is not updating filenumbers

13:23 Ks-flats 5000cts, too early, but Dave needs to close, #245-264

13:26 Close dome

13:26 Sunrise

-- JochenHeidt - 10 Nov 2013
Topic revision: r15 - 30 Mar 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback